# Julian Martinez-Iriarte, and Gabriel Montes-Rojas

### v3

@@ Abstract
Linear regressions with endogeneity are widely used to estimate causal effects. This paper studies a framework that involves two common issues: endogeneity of the regressors and heteroskedasticity that depends on endogenous regressors–i.e., endogenous heteroskedasticity. We show that the presence of endogenous heteroskedasticity in the structural regression renders the two-stage least squares estimator inconsistent. To address this issue, we propose sufficient conditions and a control function approach to identify and estimate the causal parameters of interest. We establish the limiting properties of the estimator–namely, consistency and asymptotic normality–and propose inference procedures. Monte Carlo simulations provide evidence on the finite-sample performance of the proposed methods and evaluate different implementation strategies. We revisit an empirical application on job training to illustrate the methods.
@@ 1 Introduction
This paper studies linear models with endogeneity and heteroskedasticity. At the estimation stage, endogeneity is typically addressed using instrumental variables (IV) and two-stage least squares (2SLS) procedures. Heteroskedasticity, on the other hand, is usually handled with robust inference techniques. However, when heteroskedasticity is caused by an endogenous regressor, a situation referred to as endogenous heteroskedasticity, the 2SLS estimator becomes inconsistent because the necessary exogeneity condition fails. The literature on econometric models that allow for endogenous heteroskedasticity is very limited (see, e.g., Florens, Heckman, Meghir, and Vytlacil (2008), Chen and Khan (2014), and Abrevaya and Xu (2023)), and this paper focuses specifically on that issue.
We propose alternative methods for identification, estimation, and inference for the parameter of interest in linear triangular models with endogeneity and endogenous heteroskedasticity. We study identification using a control function (CF) approach in the presence of heteroskedasticity in both the structural and first-stage equations. We first show that general nonparametric CF identification fails due to the presence of skedastic terms. We then build on Florens, Heckman, Meghir, and Vytlacil (2008) and provide conditions that address this challenge. The key restrictions for identification are a CF condition, along with a flexible parametrization of the skedastic function in the structural equation. Given a CF constructed from the first stage and the assumed structure of the skedastic function, the parameter of interest in the second stage can be identified using ordinary least squares regression.
@@
Estimation is based on the identification result and it is implemented empirically in a two-step procedure. First, the skedastic function in the first step can be estimated using flexible scale function methods suggested in Romano and Wolf (2017) and Wooldridge (2010). [^1] The parametric model in the first-stage is mainly used due to the curse of dimensionality problem, since most empirical work uses multiple control variables. In a nonparemetric case, estimation could rely on methods suggested by Jin, Su, and Xiao (2015) and Linton and Xiao (2019). Given the estimate of the skedastic function, the CF can be computed as the normalized errors from the first-stage regression. [^2] We note that under heteroskedasticity in the first-stage regression, by assuming the availability of a valid exogenous IV, the CF can be identified by using a flexible parametric model for the skedastic function. By normalizing the variance of the error term in the first-stage equation, a valid CF variable can be constructed. In the second step, one performs a simple linear regression with a CF approach. Empirically, this step uses an OLS regression of the dependent variable on the endogenous, exogenous, and CF variables augmented with their interacting terms. We establish the limiting statistical properties of the proposed estimator. Mild sufficient conditions are provided for the two-step estimator to have desired asymptotic properties, namely, consistency and asymptotic normality. In addition, we develop statistical inference procedures by estimating the asymptotic variance-covariance matrix with the generated regressors.
@@
Literature review. There is a large literature on the CF approach. See, e.g., among many others, Newey, Powell, and Vella (1999), Kim and Petrin (2022), and Han and Kaido (2025). Of particular interest are linear triangular systems is an issue of interest for econometricians together with the CF approach. Different alternatives have been studied in the literature. See the survey by Blundell and Powell (2008) and Imbens and Newey (2009) for nonparametric and semiparametric models.
The literature considering the effects of endogenous heteroskedasticity is more restricted. When treatment effects are heterogeneous among observationally identical individuals, that is, in the presence of heteroskedasticity in the structural equation, causal inference for policy evaluation is known to be more difficult. Florens, Heckman, Meghir, and Vytlacil (2008) use the control function approach to identify the average treatment effects and the effect of treatment on the treated in models with a continuous endogenous regressor whose impact is heterogeneous. Abrevaya and Xu (2023) show that when a binary treatment is endogenous and there is endogenous heteroskedasticity (EH), that is, the endogenous treatment indicator also affects the scale of the outcome variable, then standard methods for estimation of average treatment effects fail. In particular, they consider the standard IV estimator with binary treatment and show that it is an inconsistent estimator for the average treatment effect. After nonparametric identification is established, they provide closed-form estimators are provided under a linear specification for the mean and variance treatment effect, as well as the average treatment effect on the treated. In the context of local average treatment effects, Chen and Khan (2014) discuss identification and estimation of the heteroskedasticity term under different treatment statuses. As opposed to the papers listed above, we do not require particular support restrictions for the endogenous variables nor the instrumental variables. That is, we do not require continuity as in Florens, Heckman, Meghir, and Vytlacil (2008), nor discreteness as in Abrevaya and Xu (2023). [^3] Heteroskedasticity can sometimes aid in addressing endogeneity problems rather than exacerbating them, as shown in Rigobon (2003), Klein and Vella (2010), and Lewbell (2011). However, this is not the case considered in the present study.
This paper is organized as follows. Section 2 shows the inconsistency of 2SLS under endogenous heteroskedasticity. In Section 3 we propose a CF approach and discuss identification. Estimation and inference procedures are detailed in Section 4. A Monte Carlo study is provided in Section 5. Section 6 illustrates the methods with an empirical application. Finally, Section 7 concludes. Mathematical proofs are relegated to the Appendix.
2 Inconsistency of the 2SLS estimator
@@
[…]
The failure of 2SLS to identify $\alpha_{1}$ in the presence of endogenous heteroskedasticity was highlighted by Abrevaya and Xu (2023) in the case of a binary $D$. When a binary IV is available, they offer a detailed analysis of the bias and propose a novel identification and estimation strategy. In a similar context with a continuous $D$, Florens, Heckman, Meghir, and Vytlacil (2008) argue that the usual IV requirements are not sufficient for identification.
@@
Consider the following simple linear case of model (1)–(2) with $g(D)=(1+\delta D)$:
$Y$ $=D\alpha_{1}+(1+\delta D)\varepsilon$
$D$ $=m(Z,V),$
for some unspecified $m(\cdot)$ function. Here, $E[D\varepsilon]\neq 0$ through the possible correlation between $\varepsilon$ and $V$. We assume that $E[\varepsilon|Z]=E[V|Z]=0$, i.e., the instrument is valid. The 2SLS estimator of $\alpha_{1}$ is the sample counterpart of $\frac{Cov[Z,Y]}{Cov[Z,D]}$. Hence, we have that
@@
The exogeneity assumptions on $Z$ do not imply that $Cov[Z,D\varepsilon]=0$. Therefore, $\alpha_{1}$ is not identified when there is heteroskedasticity in the structural equation, i.e., $\delta\neq 0.$
The next example shows that this problem persists even if $Z\perp\!\!\!\perp(\varepsilon,V)$ as in experimental settings. To get a closed-form expression we assume a linear first stage.
Example 2.
Consider a special case of Example 1:
$Y$ $=D\alpha_{1}+(1+D\delta)\varepsilon$
$D$ $=Z\pi_{1}+(1+Z\gamma)V,$
The 2SLS estimator is the sample counterpart of $\overline{\alpha}_{1}:=\frac{E[ZY]}{E[ZD]}=\alpha_{1}+\delta\frac{E[ZV\varepsilon]}{\pi_{1}E[Z^{2}]}+\delta\gamma\frac{E[Z^{2}V\varepsilon]}{\pi_{1}E[Z^{2}]}.$ Under the stronger independence assumption: $Z\perp\!\!\!\perp(V,\varepsilon)$, the expression simplifies to
$\overline{\alpha}_{1}$ $=\alpha_{1}+\delta\gamma\frac{E[V\varepsilon]}{\pi_{1}}.$
In this case, the joint heteroskedasticity together with the endogeneity (in the form of $E[V\varepsilon]\neq 0$) is inducing the bias. Thus, in this case, one could have a situation where there is heterogeneity in the structural equation with $\delta\neq 0$, but if there is no heteroskedasticity in the first-stage, $\gamma=0$, the 2SLS is able to correctly identify the parameter of interest.
Next, we provide a formal result deriving the asymptotic bias in the 2SLS estimator under the endogenous heteroskedasticity. This not a novel results per se, but we specify it to our particular setting. We maintain the following assumptions.
@@
Assumptions 1 and 2 are standard in the literature allowing for endogeneity of $D$, the presence of observable controls $X$, as well as for the valid instruments $Z$ to be (intrinsically) exogenous.
While the instruments $Z$ are uncorrelated with $\varepsilon$ and $V$, we will see that when we consider the unobservable in (1) to be $g(D,X)\varepsilon$, then it might no longer be the case that $E[Zg(D,X)\varepsilon]=0$. That is, despite $Z$ being uncorrelated to the unobserved innovations, and being relevant, the heteroskedastic nature of the triangular system can invalidate the 2SLS procedure. As an alternative, we propose a control function approach in Section 3 below.
The next result compares the probability limit of the 2SLS estimator, denoted by $\overline{\alpha}_{1}$, to the target parameter $\alpha_{1}$. We assume that the data is generated according to the model in (1)–(2), and a researcher who is interested in $\alpha_{1}$, instruments $D$ with $Z$ in a 2SLS regression of $Y$ of $D$ and $X$ (which includes a constant). In order to deal with the presence of the additional controls $X$, we use a population version of the Frisch-Waugh-Lovell (FWL) theorem using linear projections. [^5] See Section A.2 in the Appendix for a brief review of linear projections.
@@
$\overline{\alpha}_{1}=\alpha_{1}+\Sigma_{Z}^{-1}\pi_{1}^{\prime}E[\overline{Z}g(m(X,Z,V),X)\varepsilon],$ (3)
where $\Sigma_{Z}:=\pi_{1}^{\prime}E[\overline{Z}\,\overline{Z}^{\prime}]\pi_{1}$, and $\overline{Z}^{\prime}:=Z^{\prime}-X^{\prime}E[XX^{\prime}]^{-1}E[XZ^{\prime}]$.
We provide some remarks to further understand the intuition and nature of the result in Lemma 1.
Remark 1.
If $g(\cdot)\equiv 1$, i.e., there is no heteroskedasticity in the structural equation, then there is no asymptotic bias in the 2SLS estimator because $E[\overline{Z}{g(D,X)\varepsilon}]=E[\overline{Z}{\varepsilon}]=0$ by Assumption 1.
Remark 2.
@@
Even in this simple case, the sign of the bias is not obvious.
Remark 3.
@@
Remark 4.
In this paper, we focus on the 2SLS estimator. Nevertheless, we highlight the same result is valid for a GMM estimator. The main intuition is that, under endogenous heteroskedasticity, the moment condition used for estimation would be misspecified through the function $g(\cdot)$.
3 Identification: a control function approach
In this section, we propose sufficient conditions, together with a control function approach, to identify and estimate the parameters of interest in the model given in equations (1)–(2), in the presence of both endogeneity and general heteroskedasticity.

### v4

@@ Abstract
Linear regressions with endogeneity are widely used to estimate causal effects. This paper studies a framework that involves two common practical issues: endogeneity of the regressors and heteroskedasticity that depends on endogenous regressors, i.e., endogenous heteroskedasticity. To address the inconsistency of the two-stage least squares estimator in this scenario, and recover the causal parameters of interest, we develop a framework for practical estimation and inference based on the control function approach allowing for discrete and continuous regressors. In particular, we suggest a simple two-step estimation procedure. We establish the limiting properties of the estimator, namely, consistency and asymptotic normality. In addition, we develop practical valid inference methods by proposing an estimator for the asymptotic variance-covariance matrix, and formally establishing its consistency. Monte Carlo simulations provide evidence on the finite-sample performance of the proposed methods and evaluate different implementation strategies. We revisit an empirical application on job training to illustrate the methods.
@@ 1 Introduction
This paper studies estimation of linear models with endogeneity and heteroskedasticity. At the estimation stage, endogeneity is typically addressed using instrumental variables (IV) and two-stage least squares (2SLS) procedures. Heteroskedasticity, on the other hand, is usually handled with robust inference techniques. However, when heteroskedasticity is caused by an endogenous regressor, a situation referred to as endogenous heteroskedasticity, the 2SLS estimator may become inconsistent because the necessary exogeneity condition fails. The literature on econometric models that allow for endogenous heteroskedasticity is limited (see, e.g., FlorensHeckmanMeghirVytlacil08, ChenKhan14, and AbrevayaXu23).
In this paper, we focus on this issue and develop methods for estimation and inference of parameters in linear triangular models with endogeneity and endogenous heteroskedasticity that employ a control function (CF) approach and allow for continuous and discrete regressors. Endogenous heteroskedasticity models can be viewed as a special case of nonseparable models, see, e.g., Chesher03, ChernozhukovImbensNewey07, FlorensHeckmanMeghirVytlacil08, ImbensNewey09 and Jun09. These general methods usually employ nonparametric modeling and estimation together with a CF. Differently from these papers, our interest lies in practical linear regression models in the presence of multiple covariates, such that nonparametric methods become impractical. Thus, we rely on the functional necessity of building a feasible parametric model using the CF approach. [^1] There is a large literature on the CF approach. See, e.g., among many others, NeweyPowellVella99, KimPetrin22, and han2025setvaluedcontrolfunctions. Different alternatives have been studied in the literature. See the survey by BlundellPowell08, and ImbensNewey09 for nonparametric and semiparametric models. We start by building on the general identification results in FlorensHeckmanMeghirVytlacil08 and tailor the conditions to accommodate continuous and discrete variables such that we are able to write the parameter of interest as an explicit function of observables, and thus derive the estimator of interest. The key restrictions for identification are a CF condition with a flexible parameterization, along with another parameterization of the skedastic function in the structural equation. Given a valid CF constructed from the first stage, and the structure of the skedastic function, the parameter of interest in the second stage can be identified using a standard ordinary least squares (OLS) regression. We highlight that our proposed approach includes the 2SLS as a special case when there is no endogenous heteroskedasticity and there is no correction for the first-stage heteroskedasticity, a condition that depends on the CF assumption.
@@
Estimation is implemented in a two-step procedure. First, the skedastic function in the first step can be estimated using flexible scale function methods suggested in RomanoWolf17. Given the estimate of the skedastic function, the CF can be computed as the normalized errors from the first-stage regression. In the second step, one performs a simple linear regression with a CF approach. Empirically, this step uses an OLS regression of the dependent variable on the endogenous, exogenous, and CF variables augmented with their interacting terms. We establish the limiting statistical properties of the proposed two-step estimator. Mild sufficient conditions are provided for the estimator to have desired asymptotic properties, namely, consistency and asymptotic normality. The limiting theory presented here incorporates elements of both the CF and generated regressors.
In addition, we develop practical statistical inference procedures by suggesting an estimator for the asymptotic variance-covariance matrix with the generated regressors. We formally establish the consistency of the estimator of the variance-covariance matrix. The proposed methods provide foundation for general inference procedures. Testing for general linear hypotheses is easily accommodated by Wald type tests, and non-linear hypothesis are readily available by employing the Delta method.
@@
The specific literature considering the effects of endogenous heteroskedasticity in linear models is restricted. When treatment effects are heterogeneous among observationally identical individuals, that is, in the presence of heteroskedasticity in the structural equation, causal inference for policy evaluation is known to be more difficult. FlorensHeckmanMeghirVytlacil08 use the CF approach to identify the average treatment effects and the effect of treatment on the treated in models with a continuous endogenous regressor whose impact is heterogeneous. AbrevayaXu23 show that when a binary treatment is endogenous and there is endogenous heteroskedasticity, that is, the endogenous treatment indicator also affects the scale of the outcome variable, then standard methods for estimation of average treatment effects fail. In particular, they consider the standard IV estimator with binary treatment and show that it is an inconsistent estimator for the average treatment effect. After nonparametric identification is established, estimators are provided under a linear specification for the mean and variance treatment effect, as well as the average treatment effect on the treated. In the context of local average treatment effects, ChenKhan14 discuss identification and estimation of the heteroskedasticity term under different treatment statuses. As opposed to the papers listed above, we do not require particular support restrictions for the endogenous variables nor the instrumental variables. That is, we do not require continuity as in FlorensHeckmanMeghirVytlacil08, nor discreteness as in AbrevayaXu23. Heteroskedasticity can sometimes aid in addressing endogeneity problems rather than exacerbating them, as shown in Rigobon03, KleinVella10, and Lewbell12. However, this is not the case considered in the present study.
This paper is organized as follows. Section 2 presents the model and shows the inconsistency of 2SLS under endogenous heteroskedasticity. In Section 3 we propose a CF approach and discuss identification. Estimation and inference procedures are detailed in Section 4. A Monte Carlo study is provided in Section 5. Section 6 illustrates the methods with an empirical application. Finally, Section 7 concludes. Mathematical proofs are relegated to the Appendix.
2 A linear model with endogenous heteroskedasticity
In this section, we first describe the model of interest, which contains endogenous heteroskedasticity. Second, we motivate the proposed methods by reviewing the inconsistency of the two-stage least squares in this scenario.
2.1 Model
@@
The model in equations (1) and (2) has several important features. Despite the linearity in $D$ in (1), it allows for heterogeneous effects on $Y$ due to the term $g(D,X)\varepsilon$. Because of this, the parameter of interest has an interpretation as an average effect. Indeed, if we define the potential outcome for unit $i$ when $D=d$ as $Y_{di}=d\alpha_{1}+X_{i}^{\prime}\alpha_{2}+g(d,X_{i})\varepsilon_{i}$, then $E[Y_{di}-Y_{d^{\prime}i}]=(d-d^{\prime})\alpha_{1}$. [^2] See Appendix A.1 for more about the interpretation of $\alpha_{1}$. However, the exogeneity conditions $E[\varepsilon|X,Z]=0$ and $E[V|X,Z]=0$ are not sufficient to identify $\alpha_{1}$ via 2SLS, since $E[g(D,X)\varepsilon|X,Z]$ may not be zero. Since the conditional variance of the error term in (1) is $g(D,X)^{2}Var[\varepsilon|D,X]$, we refer to it as “endogenous heteroskedasticity,” a terminology borrowed from AbrevayaXu23.
The model above is a particular case of nonadditive triangular models, where the interest lies in estimating the average linear effect, $\alpha_{1}$. This specification would allow us to study the role of the skedastic function $g$ in the 2SLS bias for $\alpha_{1}$ and specific solutions using the control function approach. A comparable model is used in AbrevayaXu23 to study the role of endogenous heteroskedasticity in treatment effects models. Furthermore, equation (2) could be rewritten as a general model $D=m(Z,X,V)$ but we use a similar skedastic structure to match equation (1). The parametric model in the first-stage is mainly used due to the curse of dimensionality problem, since most empirical work uses multiple control variables. In a nonparemetric case, estimation could rely on methods suggested by JiSuXiao15 and LintonXiao19.
2.2 Motivation: Inconsistency of the 2SLS estimator
The 2SLS is a popular method to address endogeneity in linear models. In this section we show that the 2SLS estimator estimates the average effect $\alpha_{1}$ with a bias in the presence of endogeneous heteroskedasticity. Although the result here is simple, intuitive, and similar issues were highlighted in the literature, we include this section to motivate the linear representation above.
The failure of 2SLS to identify $\alpha_{1}$ in the presence of endogenous heteroskedasticity was highlighted by AbrevayaXu23 in the case of a binary $D$. When a binary IV is available, they offer a detailed analysis of the bias and propose a novel identification and estimation strategy. In a similar context with a continuous $D$, FlorensHeckmanMeghirVytlacil08 argue that the usual exogeneity IV requirements are not sufficient for identification.
@@ 2.2 Motivation: Inconsistency of the 2SLS estimator
Consider the following simple linear case of model (1)–(2) with one endogenous variable, one instrument, $g(D)=1+\delta D$ and $h(Z)=1+Z\gamma$:
$Y$ $=D\alpha_{1}+(1+D\delta)\varepsilon$
$D$ $=Z\pi_{1}+(1+Z\gamma)V.$
Here, $E[D\varepsilon]\neq 0$ through the possible correlation between $\varepsilon$ and $V$. We assume that $E[\varepsilon|Z]=E[V|Z]=0$. The 2SLS estimator of $\alpha_{1}$ is the sample counterpart of $\frac{Cov[Z,Y]}{Cov[Z,D]}$. Hence, we have that
@@
The exogeneity assumptions on $Z$ do not imply that $Cov[Z,D\varepsilon]=0$. Therefore, $\alpha_{1}$ is not identified by the 2SLS moment conditions whenever there is heteroskedasticity in the structural equation, i.e., $\delta\neq 0.$
Under the stronger independence assumption: $Z\perp\!\!\!\perp(V,\varepsilon)$ (typical in experimental settings when $Z$ is random), the expression simplifies to
$\frac{Cov[Z,Y]}{Cov[Z,D]}$ $=\alpha_{1}+\delta\gamma\frac{E[V\varepsilon]}{\pi_{1}}$
In this case, the joint heteroskedasticity together with the endogeneity (in the form of $E[V\varepsilon]\neq 0$) is inducing the bias. Thus, in this case, one could have a situation where there is heterogeneity in the structural equation with $\delta\neq 0$, but if there is no heteroskedasticity in the first-stage, $\gamma=0$, then 2SLS is able to correctly identify the parameter of interest.
Next, we provide a formal result deriving the asymptotic bias in the 2SLS estimator under the endogenous heteroskedasticity. This not a novel results per se, but we specify it to our particular setting.
We maintain the following assumptions.
@@
Assumptions 1 and 2 are standard in the literature allowing for endogeneity of $D$, the presence of observable controls $X$, as well as for the valid instruments $Z$ to be (intrinsically) exogenous. While the instruments $Z$ are uncorrelated with $\varepsilon$ and $V$, we will see that when we consider the unobservable in (1) to be $g(D,X)\varepsilon$, then it might no longer be the case that $E[Zg(D,X)\varepsilon]=0$. That is, despite $Z$ being uncorrelated to the unobserved innovations, and being relevant, the heteroskedastic nature of the triangular system can invalidate the 2SLS procedure. As an alternative, we propose a control function approach in Section 3 below.
The next result compares the probability limit of the 2SLS estimator, denoted by $\overline{\alpha}_{1}$, to the target parameter $\alpha_{1}$. We assume that the data is generated according to the model in (1)–(2), and a researcher who is interested in $\alpha_{1}$, instruments $D$ with $Z$ in a 2SLS regression of $Y$ on $D$ and $X$ (which includes a constant). In order to deal with the presence of the additional controls $X$, we use a population version of the Frisch-Waugh-Lovell (FWL) theorem using linear projections. [^3] See Section A.2 in the Appendix for a brief review of linear projections.
@@
$\overline{\alpha}_{1}=\alpha_{1}+\Sigma_{h}^{-1}\pi_{1}^{\prime}E[\overline{Z}g(Z^{\prime}\pi_{1}+X^{\prime}\pi_{2}+h(X,Z)V,X)\varepsilon],$ (3)
where $\Sigma_{h}:=\pi_{1}^{\prime}E[\overline{Z}\,\overline{Z}^{\prime}]\pi_{1}$, and $\overline{Z}^{\prime}:=Z^{\prime}-X^{\prime}E[XX^{\prime}]^{-1}E[XZ^{\prime}]$.
We provide some further remarks to understand the intuition and nature of the result in Lemma 1.
  1. 1.
    The asymptotic bias of the 2SLS estimator in equation (3) can be succinctly written as $\Sigma_{h}^{-1}\pi_{1}^{\prime}E[\overline{Z}g(D,X)\varepsilon]$, but this might obscure the fact that $h(\cdot)$ is also playing a role through $D$, and could, in principle, amplify the bias.
  2. 2.
    If $g(\cdot)\equiv 1$, i.e., there is no heteroskedasticity in the structural equation, then there is no asymptotic bias in the 2SLS estimator because $E[\overline{Z}{g(D,X)\varepsilon}]=E[\overline{Z}{\varepsilon}]=0$ by Assumption 1. The lack of bias is regardless of $h(\cdot)$. For a general $g(\cdot)$, the form of $h(\cdot)$ may matter.
  3. 3.
    If $h(\cdot)\equiv 1$, i.e., there is no heteroskedasticity in the first-stage, then the asymptotic bias depends only on the form of $g(\cdot)$.
  4. 4.
@@
    Even in this simple case, the sign of the bias is not totally obvious.
  5. 5.
@@
  6. 6.
    In this paper, we focus on the 2SLS estimator. Nevertheless, we highlight the same result is valid for a GMM estimator. The main intuition is that, under endogenous heteroskedasticity, the moment condition used for estimation would be misspecified trough the function $g(\cdot)$.
3 A control function approach
This section discusses sufficient conditions, based on the control function (CF) approach, that allow for identification of the parameters of interest in the model given in equations (1)–(2), in the presence of both endogeneity and general heteroskedasticity. In the next section, we will suggest a practical estimator and develop inference procedures. Identification in additive separable models with endogenous heteroskedasticity using a CF strategy was initially proposed FlorensHeckmanMeghirVytlacil08. Our analysis builds on their results, with the key distinction between our approach and theirs being a trade-off between making a stochastic polynomial assumption on the CF in our approach versus their approach allowing for a general nonparametric function. The advantage of our strategy is reflected in the practicality of the estimator. We will show that, by specializing the CF to be a polynomial, the identification result shows that the parameter of interest can be written as a simple explicit function of the data, which in turn can be used for estimation and inference. Moreover, our approach allows the 2SLS-IV estimator to be nested within our model, and thus being a special case.